# CMB parity asymmetry from unitary quantum gravitational physics

### v2

@@ Appendix B Appendix B: DQFT in dS
Here the operators $d_{\pm\,\textbf{k}},\,d^{\dagger}_{\pm\,\textbf{k}}$ satisfy the similar relations like in Eq.23. The dS mode functions $\phi_{\pm\,k}=\frac{1}{\sqrt{2k}}\left(1\mp\frac{i}{k\tau}\right)e^{\mp ik\tau}$ corresponding to the Bunch-Davies vacuum of the DQFT, which recover short distance behavior of quantum fields in Minkowski Eq.22.
Appendix C DSI angular power spectrum
Following Gaztañaga and Kumar (2024), the CMB power spectrum $C_{\ell}$ in DSI is:
$C_{\ell}^{odd}=\frac{2}{9\pi}\int_{0}^{k_{c}}\frac{dk}{k}j_{\ell}^{2}\left(\frac{k}{k_{s}}\right)\mathcal{P}_{\zeta}(k)\left(1+\Delta\mathcal{P}_{v}\right)$ (27)
$C_{\ell}^{even}=\frac{2}{9\pi}\int_{0}^{k_{c}}\frac{dk}{k}j_{\ell}^{2}\left(\frac{k}{k_{s}}\right)\mathcal{P}_{\zeta}(k)\left(1-\Delta\mathcal{P}_{v}\right)$
where $\mathcal{P}_{\zeta}=A_{s}\left(\frac{k}{k_{\ast}}\right)^{n_{s}-1}$, $A_{s}=2.2\times 10^{-9}$ and
$\Delta\mathcal{P}_{v}=\left(1-n_{s}\right)\operatorname{Re}\left[\frac{2}{H_{3/2}^{(1)}\left(\frac{k}{k_{\ast}}\right)}\frac{\partial H^{(1)}_{\nu_{s}}\left(\frac{k}{k_{\ast}}\right)}{\partial\nu_{s}}\Bigg{|}_{\nu_{s}=\frac{3}{2}}\right]$ (28)
Here, $H^{(1)}_{\nu_{s}}\left(z\right),j_{\ell}\left(z\right)$ are the Hankel and Bessel functions of the first kind, $k_{s}=7\times 10^{-5}\,{\rm Mpc}^{-1}$ and $k_{c}=0.02k_{\ast}$ is the cut-scale we impose as Eq.28 is accurate enough for $-\ell$ or large angular scales. This cutoff scale is also related to the coarse-graining scale of stochastic inflation (Gaztañaga and Kumar (2024)), which determines the large wavelength modes that have already become classical on the onset of inflation while the mode $k_{\ast}$ exits the horizon. In DSI, these large-wavelength modes create parity-asymmetric inhomogeneities that correct the FLRW spacetime on large scales. The parity asymmetry ceases to exist towards the small scales because the $\mathcal{P}\mathcal{T}$ symmetry is recovered towards the short-distance scales. This can be seen in Fig. 11 of Gaztañaga and Kumar (2024). An intuitive explanation for this is that large wavelength modes experience quantum gravitational effects more than those of short wavelength modes. For the small scales $k>k_{c}$, the predictions of DSI become almost the same as with the (near) scale-invariant power spectrum as contributions from $\Delta\mathcal{P}_{v}$ term become negligible.

### v4

@@ Appendix C DQFT in dS
Here the operators $d_{(\pm)\,\textbf{k}},\,d^{\dagger}_{(\pm)\,\textbf{k}}$ satisfy the similar relations like in Eq.48. The dS mode functions are:
$\phi_{\pm\,k}=\frac{1}{\sqrt{2k}}\left(1\mp\frac{i}{k\tau}\right)e^{\mp ik\tau}=\frac{\sqrt{\mp\pi\tau}}{2}H_{3/2}^{(1)}\left(\mp k\tau\right)$ (56)
corresponding to the Bunch-Davies vacuum of the DQFT
$|0\rangle_{\rm dS}=\begin{pmatrix}|0_{+}\rangle_{\rm dS}\\ |0_{-}\rangle_{\rm dS}\end{pmatrix}$ (57)
which recovers short-distance behavior of quantum fields in Minkowski Eq. 47 i.e.,
$\hat{\phi}_{\pm}\Big{|}_{k\tau\gg 1}=\frac{1}{\sqrt{2k}}e^{\mp ik\tau}$ (58)
The two point correlations in de Sitter spacetime are identical at parity conjugate regions
$\frac{1}{a^{2}}{}_{\rm dS}\langle 0_{+}|\hat{\phi}_{+}\left(\tau,\,\textbf{x}\right)\hat{\phi}_{+}\left(\tau,\,\textbf{x}^{\prime}\right)|0_{+}\rangle_{\rm dS}=$ (59)
$\frac{1}{a^{2}}{}_{dS}\langle 0_{-}|\hat{\phi}_{-}\left(-\tau,\,-\textbf{x}\right)\hat{\varphi}_{-}\left(-\tau,\,-\textbf{x}^{\prime}\right)|0\rangle_{dS-}=\int\frac{dk}{k}\frac{\sin k\xi}{k\xi}\frac{H^{2}}{4\pi^{2}}\,,$
where $\xi=|\textbf{x}-\textbf{x}^{\prime}|$.